QUESTION 3 Suppose you are considering a project that will require an initial investment of $215,000. This project is expected to provide cash flows over the next five years as follows: $50,000, $50,000, $50,000, $50,000 and $50,000. What is the intemal rate of return for this project? At a discount rate of 13%, should you accept or reject this project?
Answer #1

answer:

Initial investment P= $215000

Cash flow C= $50000 per year

Let X be the internal rate of return.

For internal rate of return calculation

NPV=0

PV of all future cash flow-initial investment =0

PV of all future cash flow=initial investment

50000*(1-(1+X)^-5)/X=215000    PV of an annuity =C*(1-(1+r)^-n)/r

X= 5.25%    c=payment per period , r = periodic discount rate

   n=number of periods

For r=13%

NPV of project

NPV=50000*(1-(1+13%)^-5)/13% -215000=-$39138.43

Since we are getting negative NPV for this project so we don't accept this project.